### v2

@@ 3.3 Conditional beliefs
$\mathbb{P}\left(\cap_{i=1}^{n}B_{i}\cap A_{i}^{*}\right)=\mathbb{P}(B_{i})\Lambda_{i}(y_{i}^{*}),~{}i=1,\dots,n-1,~{}\mathbb{P}(A_{n}^{*})=\mathbb{P}(\cap_{i=1}^{n}B_{i})-\sum_{j=1}^{n-1}\mathbb{P}(B_{j})\Lambda_{j}(y_{j}^{*}).$
In this special case, the interplay among $\mathbb{Q}_{1},\dots,\mathbb{Q}_{n}$ is characterized by the relationship between the sets $B_{1},\dots,B_{n}$. Following the conclusion in Proposition 3, those sets $B_{1},\dots,B_{n}$ appear in the expressions of the inf-convolution and the optimal allocation, suggesting the influence of the relation of $\mathbb{Q}_{1},\dots,\mathbb{Q}_{n}$ on the inf-convolution. In particular, if the agents have a strong disagreement on the interested events such that $\mathbb{P}(\cap_{i=1}^{n}B_{i})=0$, it leads to a trivial case. Note that the crucial difference to the homogeneous case is that here, in characterising the allocations, we use the set $\left(\cap_{i=1}^{n}B_{i}\right)\cap A_{i}^{*}$ which includes both the set $A_{i}^{*}$ on which the contingent loss is paid and the conditioning set (information) on which all agents can agree. If all Lambda functions in Proposition 3 are assumed to be constants, then we immediately derive the explicit formulas for the inf-convolution of multiple $\mathrm{VaR}$ under conditional beliefs.
@@
$\mathop{\square}\displaylimits_{i=1}^{n}\mathrm{VaR}_{\alpha_{i}}^{\mathbb{Q}_{i}}(X)=\mathrm{VaR}_{1\wedge\sum_{i=1}^{n}\frac{\alpha_{i}\mathbb{P}(B_{i})}{\mathbb{P}(\cap_{i=1}^{n}B_{i})}}^{\mathbb{Q}}(X),$
@@
The results in Corollary 1 extend the elegant formula for the inf-convolution of $\mathrm{VaR}$ under homogeneous beliefs to the case with conditional beliefs; see Corollary 2 of Embrechts et al. (2018). The simplicity of the expressions for $\mathrm{VaR}$ allow some additional insight. We can see that the trivial outcome $\mathop{\square}\displaylimits_{i=1}^{n}\mathrm{VaR}_{\alpha_{i}}^{\mathbb{Q}_{i}}(X)=-\infty$ is not only achieved in the case of full disagreement between agents $\mathbb{P}(\cap_{i=1}^{n}B_{i})=0$. It also occurs under weaker disagreement, when $\sum_{i=1}^{n}\frac{\alpha_{i}\mathbb{P}(B_{i})}{\mathbb{P}(\cap_{i=1}^{n}B_{i})}\geqslant 1$. This inequality may be satisfied, when $\mathbb{P}(\cap_{i=1}^{n}B_{i})$ is sufficiently small, indicating high – but partial – disagreement in conditional beliefs, and the levels $\alpha_{i}$ are sufficiently large, indicating a high level of risk tolerance.
@@
3.4 Absolutely continuous beliefs
Next, we consider a different relation among the probability measures, absolute continuity. Due to the complexity of the problem, we here only consider the case $n=2$.
@@ 3.4 Absolutely continuous beliefs
Suppose $\mathbb{Q}_{2}<<\mathbb{Q}_{1}$ and denote $\frac{d\mathbb{Q}_{2}}{d\mathbb{Q}_{1}}$ by $\eta$. For $\Lambda_{i}\in\mathcal{H}$ with $10<\lambda_{i}^{-}\leqslant\lambda_{i}^{+}<$ 1, we have
@@
where $g_{x}(t)=\mathbb{E}^{\mathbb{Q}_{1}}(\eta\mathds{1}_{\{X>x,U_{\eta}^{\mathbb{Q}_{1}}\leqslant x_{t}\}}),~{}t\in[0,1-F_{X}^{\mathbb{Q}_{1}}(x)]$, with $x_{t}$ satisfying $\mathbb{Q}_{1}(X>x,U_{\eta}^{\mathbb{Q}_{1}}\leqslant x_{t})=t$. Moreover, if $x^{*}=\Lambda_{1}\mathrm{VaR}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\Lambda_{2}\mathrm{VaR}^{\mathbb{Q}_{2}}(X)>-\infty$ and $g_{x^{*}}((1-F_{X}^{\mathbb{Q}_{1}}(x^{*})-\Lambda_{1}(y^{*}))_{+})\leqslant\Lambda_{2}(x^{*}-y^{*})$ for some $y^{*}\in\mathbb{R}$, then the optimal allocation is given by
@@
where $A^{*}=\{X>x^{*},U_{\eta}^{\mathbb{Q}_{1}}\leqslant z^{*}\}$ with $z^{*}$ satisfying $\mathbb{Q}_{1}(X>x^{*},U_{\eta}^{\mathbb{Q}_{1}}\leqslant z^{*})=(1-F_{X}^{\mathbb{Q}_{1}}(x^{*})-\Lambda_{1}(y^{*}))_{+}$.
@@
    If $X$ and $\eta$ are comonotonic, then $g_{x}(t)=\int_{F_{X}^{\mathbb{Q}_{1}}(x)}^{F_{X}^{\mathbb{Q}_{1}}(x)+t}F_{\eta}^{\mathbb{Q}_{1},-1}(s)\mathrm{d}s,~{}t\in[0,1-F_{X}^{\mathbb{Q}_{1}}(x)]$ and $A^{*}=\{X>x^{*},U_{\eta}^{\mathbb{Q}_{1}}\leqslant(1-\Lambda_{1}(y^{*}))\vee F_{X}^{\mathbb{Q}_{1}}(x^{*})\}$;
@@
    If $X$ and $\eta$ are countermonotonic, then $g_{x}(t)=\int_{0}^{t}F_{\eta}^{\mathbb{Q}_{1},-1}(s)\mathrm{d}s,~{}t\in[0,1-F_{X}^{\mathbb{Q}_{1}}(x)]$ and $A^{*}=\{X>x^{*},U_{\eta}^{\mathbb{Q}_{1}}\leqslant(1-F_{X}^{\mathbb{Q}_{1}}(x^{*})-\Lambda_{1}(y^{*}))_{+}\}$;
@@
    If $X$ and $\eta$ are independent under $\mathbb{Q}_{1}$, then $g_{x}(t)=(1-F_{X}^{\mathbb{Q}_{1}}(x))\int_{0}^{t/(1-F_{X}^{\mathbb{Q}_{1}}(x))}F_{\eta}^{\mathbb{Q}_{1},-1}(s)\mathrm{d}s,~{}t\in[0,1-F_{X}^{\mathbb{Q}_{1}}(x)]$ with the convention that $/0=0$ 0 and $A^{*}=\{X>x^{*},U_{\eta}^{\mathbb{Q}_{1}}\leqslant\frac{(1-F_{X}^{\mathbb{Q}_{1}}(x^{*})-\Lambda_{1}(y^{*}))_{+}}{1-F_{X}^{\mathbb{Q}_{1}}(x^{*})}\}$ with $U_{\eta}^{\mathbb{Q}_{1}}$ being independent of $X$ under $\mathbb{Q}_{1}$.
Note that the optimal set $A^{*}$ in (i) and (ii) have very different structures. To demonstrate this, suppose both $X$ and $\eta$ are continuous random variables under $\mathbb{Q}_{1}$. In (i), $A^{*}$ can be rewritten as $A^{*}=\{F_{X}^{\mathbb{Q}_{1}}(x^{*})<U_{X}^{\mathbb{Q}_{1}}<(1-\Lambda_{1}(y^{*}))\vee F_{X}^{\mathbb{Q}_{1}}(x^{*})\}$. Hence, the contingent loss of agent 2 takes place on a set that does not include the full right tail of $X$. This is understandable because the comonotonicity of $X$ and $\eta$ means that agent 2 considers the distribution of aggregate loss more risky in first-order stochastic dominance, $F_{X}^{\mathbb{Q}_{2}}\succeq_{\mathrm{st}}F_{X}^{\mathbb{Q}_{1}}$ (see e.g. Pesenti et al. (2019)). In the alternative case (ii) of comonotonicity, we can write $A^{*}=\{U_{X}^{\mathbb{Q}_{1}}>F_{X}^{\mathbb{Q}_{1}}(x^{*})+\Lambda_{1}(y^{*})\}$, such that agent 2 is exposed to the full tail of $X$, consistent with $F_{X}^{\mathbb{Q}_{1}}\succeq_{\mathrm{st}}F_{X}^{\mathbb{Q}_{2}}$. Hence, by (4), we see that if $X$ and $\eta$ are comonotonic, a lower contingent loss is allocated to the second agent; if $X$ and $\eta$ are countermonotonic, a higher contingent loss is allocated to the second agent.
@@
Suppose $\mathbb{Q}_{2}<<\mathbb{Q}_{1}$ and denote $\frac{d\mathbb{Q}_{2}}{d\mathbb{Q}_{1}}$ by $\eta$. For $\alpha_{1},\alpha_{2}\in(0,1)$, we have
@@
where $g_{x}(t)=\mathbb{E}^{\mathbb{Q}_{1}}(\eta\mathds{1}_{\{X>x,U_{\eta}^{\mathbb{Q}_{1}}\leqslant x_{t}\}}),~{}t\in[0,1-F_{X}^{\mathbb{Q}_{1}}(x)]$, with $x_{t}$ satisfying $\mathbb{Q}_{1}(X>x,U_{\eta}^{\mathbb{Q}_{1}}\leqslant x_{t})=t$.
@@
    with the convention that $/0=0$ 0.
4 Inf-convolution of $\Lambda\mathrm{VaR}$ and one monotone risk measure with heterogenous beliefs
Throughout this section, we set $\mathcal{X}=L^{\infty}$. We investigate the risk sharing problem between two agents, considering the inf-convolution of $\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}$ and a monotone risk measure $\rho^{\mathbb{Q}_{2}}$, where $\mathbb{Q}_{2}$ may be different from $\mathbb{Q}_{1}$. Note that $\rho$ here is very general including many commonly used risk measures as special cases.
@@ 4 Inf-convolution of $\Lambda\mathrm{VaR}$ and one monotone risk measure with heterogenous beliefs
Suppose $\Lambda\in\mathcal{H}$ with $10<\lambda^{-}\leqslant\lambda^{+}<$ 1 and $\rho$ is monotone. Then we have
@@
Our conclusion in Theorem 5 extends the results of Embrechts et al. (2020) for $n=2$, where the inf-convolution of $\mathrm{VaR}$ or $\mathrm{ES}$ with heterogeneous beliefs is considered. If $\mathbb{Q}_{1}=\mathbb{Q}_{2}$, then Theorem 5 coincides with the results in Theorem 3 of Liu (2024). We notice that for the scenario that $\mathbb{Q}_{1}=\mathbb{Q}_{2}$ and $\Lambda$ is a constant, Theorem 5 extends the results in Theorem 2 of Liu et al. (2022) ($\rho$ is there considered to be a monetary tail risk measure) and Theorem 5.3 of Wang and Wei (2018) ($\rho$ is there a distortion risk measure). The result of Theorem 5 is still new even if $\Lambda\mathrm{VaR}$ is reduced to $\mathrm{VaR}$. Let us now consider the case $\mathbb{Q}_{i}(\cdot)=\mathbb{P}(\cdot|B_{i})$ for some $B_{i}\in\mathcal{F}$ with $\mathbb{P}(B_{i})>0$. Under this setup, the contingent set $B$ in (5) is given explicitly.
@@
Suppose $\mathbb{Q}_{i}(\cdot)=\mathbb{P}(\cdot|B_{i})$ for some $B_{i}\subset\Omega$ with $\mathbb{P}(B_{i})>0,~{}i=1,2$. Moreover, suppose $\Lambda\in\mathcal{H}$ with $10<\lambda^{-}\leqslant\lambda^{+}<$ 1 and $\rho$ is monotone. Then we have
@@
Let $\mathcal{G}=\{g:[0,1]\to[0,1]|~{}g~{}\text{is increasing and left-continuous satisfying}~{}g(0)=0~{}\text{and}~{}g(1)=1\}$. For $g\in\mathcal{H}$, the distortion risk measure $\rho_{g}^{\mathbb{Q}}$ is defined as
@@
Furthermore, for an increasing function $u$, the expected utility is defined as $\mathbb{E}_{u}^{\mathbb{Q}}:X\mapsto\mathbb{E}^{\mathbb{Q}}(u(X))$.
In the following corollary, let $\mathbb{Q}_{i}(\cdot)=\mathbb{P}(\cdot|B_{i})$ for some $B_{i}\subset\Omega$ with $\mathbb{P}(B_{1}\cap B_{2})>0$, indicating that beliefs are not totally incompatible, and $\mathbb{Q}(\cdot)=\mathbb{P}(\cdot|B_{1}\cap B_{2})$, which in a sense represents the consensus view of the two agents. Applying Proposition 5, we obtain the inf-convolution of $\mathrm{\Lambda\mathrm{VaR}}$ and $\rho_{g}/\mathbb{E}_{u}/\Lambda_{1}\mathrm{VaR}^{+}$ under heterogeneous beliefs, respectively. In order to simplify the expression, we suppose $\lambda^{+}$ is attainable for $\Lambda$, i.e., there is $x\in\mathbb{R}$ such that $\Lambda(x)=\lambda^{+}$.
@@
Suppose $\Lambda\in\mathcal{H}$ with $10<\lambda^{-}\leqslant\lambda^{+}<$ 1 and $\lambda^{+}$ is attainable. Then the following conclusion holds with $\beta=\frac{(1-\lambda^{+})\mathbb{P}(B_{1})-\mathbb{P}(B_{1}\cap B_{2}^{c})}{\mathbb{P}(B_{2})}$.
@@
    For $g\in\mathcal{G}$, if $\mathbb{Q}_{1}(B_{2})\leqslant\lambda^{+}$, then $\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\rho_{g}^{\mathbb{Q}_{2}}(X)=-\infty$; if $\mathbb{Q}_{1}(B_{2})>\lambda^{+}$ then

### v3

@@ 3.3 Conditional beliefs
$\mathbb{P}\left((\cap_{k=1}^{n}B_{k})\cap A_{i}^{*}\right)=\mathbb{P}(B_{i})\Lambda_{i}(y_{i}^{*}),~i=1,\dots,n-1,~\mathbb{P}(A_{n}^{*})=\mathbb{P}(\cap_{k=1}^{n}B_{k})-\sum_{j=1}^{n-1}\mathbb{P}(B_{j})\Lambda_{j}(y_{j}^{*}).$
In this special case, the interplay among $\mathbb{Q}_{1},\dots,\mathbb{Q}_{n}$ is characterized by the relationship between the sets $B_{1},\dots,B_{n}$. Following the conclusion in Proposition 3, those sets $B_{1},\dots,B_{n}$ appear in the expressions of the inf-convolution and the optimal allocation, suggesting the influence of the relation of $\mathbb{Q}_{1},\dots,\mathbb{Q}_{n}$ on the inf-convolution. In particular, if the agents have a strong disagreement on the interested events such that $\mathbb{P}(\cap_{i=1}^{n}B_{i})=0$, it leads to a trivial case. Note that the crucial difference to the homogeneous case is that here, in characterising the allocations, we use the set $\left(\cap_{k=1}^{n}B_{k}\right)\cap A_{i}^{*}$ which includes both the set $A_{i}^{*}$ on which the contingent loss is paid and the conditioning set (information) on which all agents can agree. If all Lambda functions in Proposition 3 are assumed to be constants, then we immediately derive the explicit formulas for the inf-convolution of multiple $\mathrm{VaR}$ under conditional beliefs.
@@
$\mathop{\square}\displaylimits_{i=1}^{n}\mathrm{VaR}_{\alpha_{i}}^{\mathbb{Q}_{i}}(X)=\mathrm{VaR}_{1\wedge\sum_{i=1}^{n}\frac{\alpha_{i}\mathbb{P}(B_{i})}{\mathbb{P}(\cap_{j=1}^{n}B_{j})}}^{\mathbb{Q}}(X),$
@@
The results in Corollary 1 extend the elegant formula for the inf-convolution of $\mathrm{VaR}$ under homogeneous beliefs to the case with conditional beliefs; see Corollary 2 of Embrechts et al. (2018). The simplicity of the expressions for $\mathrm{VaR}$ allow some additional insight. We can see that the trivial outcome $\mathop{\square}\displaylimits_{i=1}^{n}\mathrm{VaR}_{\alpha_{i}}^{\mathbb{Q}_{i}}(X)=$-$\infty$ is not only achieved in the case of full disagreement between agents $\mathbb{P}(\cap_{i=1}^{n}B_{i})=0$. It also occurs under weaker disagreement, when $\sum_{i=1}^{n}\frac{\alpha_{i}\mathbb{P}(B_{i})}{\mathbb{P}(\cap_{j=1}^{n}B_{j})}\geqslant 1$. This inequality may be satisfied, when $\mathbb{P}(\cap_{j=1}^{n}B_{j})$ is sufficiently small, indicating high – but partial – disagreement in conditional beliefs, and the levels $\alpha_{i}$ are sufficiently large, indicating a high level of risk tolerance.
@@
Remark.
The law-invariance of the inf-convolution for some risk functionals was discussed in Liu et al. (2020), showing that the law-invariance of the inf-convolution under homogeneous beliefs is guaranteed by some mild condition on the risk functionals if they are law-invariant. Proposition 4 shows that the inf-convolution of $\Lambda\mathrm{VaR}$ under homogeneous beliefs is law-invariant. However, if the beliefs are heterogeneous, the conclusion may not hold, which can be seen from Corollary 1. Suppose $\mathbb{Q}_{i}(\cdot)=\mathbb{P}(\cdot|B_{i})$ for some $B_{i}\subset\Omega$ with $\mathbb{P}(\cap_{i=1}^{n}B_{i})>0$ and $\alpha_{i}\in(0,1)$ satisfying $\sum_{i=1}^{n}\frac{\alpha_{i}\mathbb{P}(B_{i})}{\mathbb{P}(\cap_{j=1}^{n}B_{j})}<1$. By Corollary 1, we have
$\mathop{\square}\displaylimits_{i=1}^{n}\mathrm{VaR}_{\alpha_{i}}^{\mathbb{Q}_{i}}(X)=\mathrm{VaR}_{\sum_{i=1}^{n}\frac{\alpha_{i}\mathbb{P}(B_{i})}{\mathbb{P}(\cap_{j=1}^{n}B_{j})}}^{\mathbb{Q}}(X),$
where $\mathbb{Q}(\cdot)=\mathbb{P}(\cdot|\cap_{i=1}^{n}B_{i})$. One can easily check that $X\overset{\mathbb{Q}_{i}}{=}Y$ for all $i=1,\dots,n$ does not imply $X\overset{\mathbb{Q}}{=}Y$. Hence, $\mathop{\square}\displaylimits_{i=1}^{n}\mathrm{VaR}_{\alpha_{i}}^{\mathbb{Q}_{i}}(X)\neq\mathop{\square}\displaylimits_{i=1}^{n}\mathrm{VaR}_{\alpha_{i}}^{\mathbb{Q}_{i}}(Y)$ may hold.
3.4 General beliefs with $n=2$
Next, we consider the case $n=2$. By Theorem A.17 of Föllmer and Schied (2016), for probability measures $\mathbb{Q}_{1}$ and $\mathbb{Q}_{2}$, we have
$\mathbb{Q}_{2}(B)=\mathbb{E}^{\mathbb{Q}_{1}}(\eta\mathds{1}_{B})+\mathbb{Q}_{2}(B\cap N)~\text{for all}~B\in\mathcal{F},$ (4)
where $N\in\mathcal{F}$ satisfying $\mathbb{Q}_{1}(N)=0$ and $\eta=\frac{d\mathbb{Q}_{2}}{d\mathbb{Q}_{1}}\mathds{1}_{N^{c}}$. If $\mathbb{Q}_{2}(N)=0$, then $\mathbb{Q}_{2}<<\mathbb{Q}_{1}$ on $\Omega$; if $\mathbb{Q}_{2}(N)>0$, then $\mathbb{Q}_{2}<<\mathbb{Q}_{1}$ on $N^{c}$ but not on $N$. In this subsection, the relation of $\mathbb{Q}_{1}$ and $\mathbb{Q}_{2}$ is represented by (4).
@@ 3.4 General beliefs with $n=2$
For $\Lambda_{i}\in\mathcal{H}$ with $0<\lambda_{i}^{-}\leqslant\lambda_{i}^{+}<1$, we have
@@
where
$g_{x}(t)=\mathbb{E}^{\mathbb{Q}_{1}}(\eta\mathds{1}_{\{X>x\}\cap\{U_{\eta}^{\mathbb{Q}_{1}}\leqslant x_{t}\}}),~t\in[0,1-F_{X}^{\mathbb{Q}_{1}}(x)],$
with $x_{t}$ satisfying $\mathbb{Q}_{1}(\{X>x\}\cap\{U_{\eta}^{\mathbb{Q}_{1}}\leqslant x_{t}\})=t$. Moreover, if $x^{*}=\Lambda_{1}\mathrm{VaR}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\Lambda_{2}\mathrm{VaR}^{\mathbb{Q}_{2}}(X)>$-$\infty$ and $g_{x^{*}}((1-F_{X}^{\mathbb{Q}_{1}}(x^{*})-\Lambda_{1}(y^{*}))_{+})\leqslant\Lambda_{2}(x^{*}-y^{*})$ for some $y^{*}\in\mathbb{R}$, then the optimal allocation is given by
@@
where $A^{*}=\{X>x^{*}\}\cap\{U_{\eta}^{\mathbb{Q}_{1}}\leqslant z^{*}\}\cap N^{c}$ with $z^{*}$ satisfying $\mathbb{Q}_{1}(\{X>x^{*}\}\cap\{U_{\eta}^{\mathbb{Q}_{1}}\leqslant z^{*}\})=(1-F_{X}^{\mathbb{Q}_{1}}(x^{*})-\Lambda_{1}(y^{*}))_{+}$.
An immediate conclusion from Theorem 4 is the following result, showing that the inf-convolution is -$\infty$ if $\mathbb{Q}_{2}(N)$ is large enough.
Corollary 2.
Under the assumption of Theorem 4, if $\mathbb{Q}_{2}(N)\geqslant 1-\lambda_{2}^{-}$, then
$\Lambda_{1}\mathrm{VaR}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\Lambda_{2}\mathrm{VaR}^{\mathbb{Q}_{2}}(X)=-\infty.$
Note that the idea used in the proof of Theorem 4 is not valid for the case $n\geqslant 3$. For the case $n\geqslant 3$, it may require a new method to simplify the main result obtained in Theorem 1.
@@
    If $X$ and $\eta$ are comonotonic, then $g_{x}(t)=\int_{F_{X}^{\mathbb{Q}_{1}}(x)}^{F_{X}^{\mathbb{Q}_{1}}(x)+t}F_{\eta}^{\mathbb{Q}_{1},-1}(s)\mathrm{d}s,~t\in[0,1-F_{X}^{\mathbb{Q}_{1}}(x)]$ and $A^{*}=\{X>x^{*}\}\cap\{U_{\eta}^{\mathbb{Q}_{1}}\leqslant(1-\Lambda_{1}(y^{*}))\vee F_{X}^{\mathbb{Q}_{1}}(x^{*})\}\cap N^{c}$;
@@
    If $X$ and $\eta$ are countermonotonic, then $g_{x}(t)=\int_{0}^{t}F_{\eta}^{\mathbb{Q}_{1},-1}(s)\mathrm{d}s,~t\in[0,1-F_{X}^{\mathbb{Q}_{1}}(x)]$ and $A^{*}=\{X>x^{*}\}\cap\{U_{\eta}^{\mathbb{Q}_{1}}\leqslant(1-F_{X}^{\mathbb{Q}_{1}}(x^{*})-\Lambda_{1}(y^{*}))_{+}\}\cap N^{c}$;
@@
    If $X$ and $\eta$ are independent under $\mathbb{Q}_{1}$, then $g_{x}(t)=(1-F_{X}^{\mathbb{Q}_{1}}(x))\int_{0}^{t/(1-F_{X}^{\mathbb{Q}_{1}}(x))}F_{\eta}^{\mathbb{Q}_{1},-1}(s)\mathrm{d}s,~t\in[0,1-F_{X}^{\mathbb{Q}_{1}}(x)]$ with the convention that $0/0=0$ and
    $A^{*}=\{X>x^{*}\}\cap\left\{U_{\eta}^{\mathbb{Q}_{1}}\leqslant\frac{(1-F_{X}^{\mathbb{Q}_{1}}(x^{*})-\Lambda_{1}(y^{*}))_{+}}{1-F_{X}^{\mathbb{Q}_{1}}(x^{*})}\right\}\cap N^{c}$
    with $U_{\eta}^{\mathbb{Q}_{1}}$ being independent of $X$ under $\mathbb{Q}_{1}$.
Note that the optimal set $A^{*}$ in (i) and (ii) have very different structures. To demonstrate this, suppose both $X$ and $\eta$ are continuous random variables under $\mathbb{Q}_{1}$ and $N=\emptyset$. In (i), $A^{*}$ can be rewritten as $A^{*}=\{F_{X}^{\mathbb{Q}_{1}}(x^{*})<U_{X}^{\mathbb{Q}_{1}}<(1-\Lambda_{1}(y^{*}))\vee F_{X}^{\mathbb{Q}_{1}}(x^{*})\}$. Hence, the contingent loss of agent 2 takes place on a set that does not include the full right tail of $X$. This is understandable because the comonotonicity of $X$ and $\eta$ means that agent 2 considers the distribution of aggregate loss more risky in first-order stochastic dominance, $F_{X}^{\mathbb{Q}_{2}}\succeq_{\mathrm{st}}F_{X}^{\mathbb{Q}_{1}}$ (see e.g. Pesenti et al. (2019)). In the alternative case (ii) of comonotonicity, we can write $A^{*}=\{U_{X}^{\mathbb{Q}_{1}}>F_{X}^{\mathbb{Q}_{1}}(x^{*})+\Lambda_{1}(y^{*})\}$, such that agent 2 is exposed to the full tail of $X$, consistent with $F_{X}^{\mathbb{Q}_{1}}\succeq_{\mathrm{st}}F_{X}^{\mathbb{Q}_{2}}$. Hence, by (5), we see that if $X$ and $\eta$ are comonotonic, a lower contingent loss is allocated to the second agent; if $X$ and $\eta$ are countermonotonic, a higher contingent loss is allocated to the second agent.
@@
For $\alpha_{1},\alpha_{2}\in(0,1)$, we have
@@
where $g_{x}(t)=\mathbb{E}^{\mathbb{Q}_{1}}(\eta\mathds{1}_{\{X>x,U_{\eta}^{\mathbb{Q}_{1}}\leqslant x_{t}\}}),~t\in[0,1-F_{X}^{\mathbb{Q}_{1}}(x)]$, with $x_{t}$ satisfying $\mathbb{Q}_{1}(\{X>x\}\cap\{U_{\eta}^{\mathbb{Q}_{1}}\leqslant x_{t}\})=t$.
@@
    with the convention that $0/0=0$.
4 Inf-convolution of $\Lambda\mathrm{VaR}$ and one monotone risk measure with heterogeneous beliefs
In this section, we investigate the risk sharing problem between two agents, considering the inf-convolution of $\Lambda\mathrm{VaR}$ and a monotone risk measure $\rho$ under the probability measures $\mathbb{Q}_{1}$ and $\mathbb{Q}_{2}$, where $\mathbb{Q}_{2}$ may be different from $\mathbb{Q}_{1}$. Note that $\rho$ here is very general including many commonly used risk measures as special cases. In our results below, we consider both the cases $\mathcal{X}=L^{\infty}$ and unbounded $\mathcal{X}$, each with different assumptions on $\rho$.
@@ 4 Inf-convolution of $\Lambda\mathrm{VaR}$ and one monotone risk measure with heterogeneous beliefs
Suppose $\Lambda\in\mathcal{H}$ with $0<\lambda^{-}\leqslant\lambda^{+}<1$ and $\rho$ is monotone. If either (i) $\mathcal{X}=L^{\infty}$ and $\rho$ is law-invariant under $\mathbb{Q}_{2}$ or (ii) $\mathcal{X}$ is unbounded and $\rho$ is an $\varepsilon$-tail risk measure under $\mathbb{Q}_{2}$ for some $\varepsilon\in(0,1)$, then we have
@@
Our conclusion in Theorem 5 extends the results of Embrechts et al. (2020) for $n=2$, where the inf-convolution of $\mathrm{VaR}$ or $\mathrm{ES}$ with heterogeneous beliefs is considered. If $\mathbb{Q}_{1}=\mathbb{Q}_{2}$, then Theorem 5 coincides with the results in Theorem 3 of Liu (2025). We notice that for the scenario that $\mathbb{Q}_{1}=\mathbb{Q}_{2}$ and $\Lambda$ is a constant, Theorem 5 extends the results in Theorem 2 of Liu et al. (2022) ($\rho$ is there considered to be a monetary tail risk measure) and Theorem 5.3 of Wang and Wei (2018) ($\rho$ is there a distortion risk measure). The result of Theorem 5 is still new even if $\Lambda\mathrm{VaR}$ is reduced to $\mathrm{VaR}$. Let us now consider the case $\mathbb{Q}_{i}(\cdot)=\mathbb{P}(\cdot|B_{i})$ for some $B_{i}\in\mathcal{F}$ with $\mathbb{P}(B_{i})>0$, where $\mathbb{P}$ is atomless. Under this setup, the contingent set $B$ in (6) is given explicitly.
@@
Suppose $\mathbb{Q}_{i}(\cdot)=\mathbb{P}(\cdot|B_{i})$ for some $B_{i}\subset\Omega$ with $\mathbb{P}(B_{i})>0,~i=1,2$. Moreover, suppose $\Lambda\in\mathcal{H}$ with $0<\lambda^{-}\leqslant\lambda^{+}<1$ and $\rho$ is monotone. If either (i) $\mathcal{X}=L^{\infty}$ and $\rho$ is law-invariant under $\mathbb{Q}_{2}$ or (ii) $\mathcal{X}$ is unbounded and $\rho$ is an $\varepsilon$-tail risk measure under $\mathbb{Q}_{2}$ for some $\varepsilon\in(0,1)$, then we have
@@
Let $\mathcal{G}=\{g:[0,1]\to[0,1]|~g~\text{is increasing and left-continuous satisfying}~g(0)=0~\text{and}~g(1)=1\}$. For $g\in\mathcal{H}$, the distortion risk measure $\rho_{g}$ under $\mathbb{Q}$ is defined as
@@
In the following corollary, let $\mathbb{Q}_{i}(\cdot)=\mathbb{P}(\cdot|B_{i})$ for some $B_{i}\subset\Omega$ with $\mathbb{P}(B_{1}\cap B_{2})>0$, indicating that beliefs are not totally incompatible, and $\mathbb{Q}(\cdot)=\mathbb{P}(\cdot|B_{1}\cap B_{2})$, which in a sense represents the consensus view of the two agents. Applying Proposition 5, we obtain the inf-convolution of $\mathrm{\Lambda\mathrm{VaR}}$ and $\rho_{g}/\Lambda_{1}\mathrm{VaR}^{+}$ under heterogeneous beliefs, respectively. In order to simplify the expression, we suppose $\lambda^{+}$ is attainable for $\Lambda$, i.e., there is $x\in\mathbb{R}$ such that $\Lambda(x)=\lambda^{+}$.
@@
Suppose $\Lambda\in\mathcal{H}$ with $0<\lambda^{-}\leqslant\lambda^{+}<1$ and $\lambda^{+}$ is attainable. Then the following conclusion holds with $\beta=\frac{(1-\lambda^{+})\mathbb{P}(B_{1})-\mathbb{P}(B_{1}\cap B_{2}^{c})}{\mathbb{P}(B_{2})}$.
@@
    For $g\in\mathcal{G}$ satisfying $g(\varepsilon)=1$ for some $\varepsilon\in(0,1)$, if $\mathbb{Q}_{1}(B_{2})\leqslant\lambda^{+}$, then $\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\rho_{g}^{\mathbb{Q}_{2}}(X)=$-$\infty$; if $\mathbb{Q}_{1}(B_{2})>\lambda^{+}$ then